### v1

Abstract
An enhancement in seismic measuring instrumentation has been proven to have implications in the quantity of observed earthquakes, since denser networks usually allow recording more events. However, phenomena such as strong earthquakes or even aseismic transients, as slow slip earthquakes, may alter the occurrence of earthquakes. In the field of seismology, it is a standard practice to model background seismicity as a Poisson process. Based on this idea, this work proposes a model that can incorporate the evolving spatial intensity of Poisson processes over time (i.e., we include temporal changes in the background seismicity when modeling). In recent years, novel methodologies have been developed for quantifying the uncertainty in the estimation of the background seismicity in homogeneous cases using Bayesian non-parametric techniques. This work proposes a novel methodology based on graphical Dirichlet processes for incorporating spatial and temporal inhomogeneities in background seismicity. The proposed model in this work is applied to study the seismicity in the southern Mexico, using recorded data from 2000 to 2015.
keywords:
ETAS, Dirichlet Process, Bayesian Nonparametrics, Background seismicity
1 Introduction
Since its introduction, the Epidemic Type Aftershock Sequence (ETAS) model [ogata1988statistical] has proved to be one of the most important tools for assessing stochastic seismicity, in which earthquake activity is modeled using Hawkes’ process [Hawkes1971Spectra]. In this context, the work by [ogata1988statistical] assumed that the mainshocks (or background earthquakes) originate from a Poisson process.
Recently, due to the importance of having uncertainty quantification of the estimators, such authors as [ross2022semiparametric] and [molkenthin2022gp] have worked with Bayesian non-parametric approaches to model a temporally homogeneous background intensity function $\mu(x,y)$ in equation (1) using Dirichlet processes mixtures (DPM) [antoniak1974mixtures] and Gaussian processes, respectively.
Since the homogeneity of the ETAS model can be affected by tectonic or pore pressure changes [hainzl2005detecting, hainzl2013impact, marsan2013monitoring], such phenomena may have implications for the occurrence of tremors. However, these changes can be produced by strong earthquakes [cruz2021short], and can also lead to an increase in the seismicity rate [nishikawa2017detection].
Furthermore, the presence of slow slip events—slow fault slip phenomena that mainly occur along the plate interface—can alter seismic activity [marsan2013slow, fukuda2018variability, nishikawa2023development], as well as periods of microearthquake quiescence [matthews1988statistical]. Moreover, not only variations in seismicity can affect ETAS model estimates, but also factors such as short-term aftershock incompleteness [hainzl2016apparent, asayesh2025improved] and changes in the cut-off magnitude [hainzl2022etas]. These temporary changes in activity emphasize the need for non-homogeneous models over time for the background activity function, such as the one introduced in this paper.
Consequently, this work contribution lies on the following key points:
  • We propose a nonparametric Bayesian approach to estimate the background seismic function based on Graphical Dirichlet Processes (GDP) [chakrabarti2024graphical], which allows for inhomogeneities in both space and time.
  • We focus on inferring the background intensity adding inhomogeneities in time, i.e. $\mu(x,y,t)$, following a GDP as proposed by [chakrabarti2024graphical]. In the context of tectonic activity, this variation of the background seismic activity model holds significance, as it may provide valuable insights into the actual tectonic physics that govern earthquake occurrence.
  • We only study the background intensity function because the [sawires2019updated] database is used. Its main feature lies in the fact that aftershocks, foreshocks, and earthquake swarms were already removed using the methodology proposed by [gardner1974sequence], leaving only mainshocks (background events) available (i.e., we will only consider the Poissonian intensity).
As discussed in detail in Section 2, the usage of hidden variables in our methodology can be extended and incorporated in the context of the ETAS model, enhancing the methodology by [ross2022semiparametric].
In the frequentist context, [li2020space] have used piecewise constant functions and kernel density estimators to model an inhomogeneous space-time background activity function in the context of the Hawkes process. They assumed that
$\mu(x,y,t)=\gamma f(x,y)v(t),$
where $\gamma\in\mathbb{R}$, $f$ is a probability density function, and $v$ is a real-valued function. Nevertheless, the assumption that space changes can be separated from changes in time is generally not realistic for earthquake activity.
In the non-parametric Bayesian framework, the estimation of $\mu(x,y,t)$ can be done using the Generalized Spatial Dirichlet Process introduced by [duan2007generalized]. Their approach allows modeling different probability density functions across space and discretized time, where the weights and point masses in the stick-breaking representation [sethuraman1994constructive] are assumed to vary smoothly over space and time. However, their formulation requires modeling the weights through an auxiliary thresholding scheme based on a collection of independent stationary Gaussian random fields. Since we keep the mass points fixed in our work, we have chosen to use the representation proposed by [chakrabarti2024graphical], which allows the weights to vary through a hierarchical model with beta distributions.
It is also worth mentioning the [fonseca2017dynamic] work, which deals with the spatio-temporal problem that occurs when point data are not directly observed and only aggregated counts are available. In this context, [kottas2007bayesian] proposed a non-parametric Bayesian approach to estimate the time-homogeneous spatial Poisson-process intensity function using Dirichlet processes, work that has been extended to the non-homogeneous case by [xiao2015modeling] using dependent Dirichlet processes (DDP) [maceachern1999dependent]. If the reader is interested in DDP and their stick-breaking representation, we suggest consulting the work [quintana2022dependent].
The main difference between the methodology by [xiao2015modeling] and the one proposed in our work lies on the DDP structure; while Kottas keeps the weights of the mixing elements fixed over time during the Dirichlet process (while varying their locations), we keep the locations fixed and move the weights, which will be discussed in detail in section 2.
The idea of keeping the mixture locations constant over time is reasonable, given that regions where tectonic plates interact require very long periods to undergo significant changes. Therefore, the main seismogenic regions will remain stable in their locations. When a change in seismicity occurs, for example, due to a strong earthquake or a slow slip event (SSE), the weight associated with the mixture component in these regions will deviate from zero, but only during the periods near the occurrence of the event. This, combined with the fact that the number of parameters to be estimated using the GDP is smaller, makes the approach presented in this work particularly attractive in the context of seismology.
2 Methodology
As is pointed out by [Fox2016Spatially], the usual approach to model the mainshocks (i.e., earthquakes that are not produced by other earthquakes, as the aftershocks), is to use an inhomogeneous Poisson Process in space and constant over time with intensity $\mu(x,y)$, and the rate of aftershocks is modeled using a triggering function $\nu$ depending on the magnitude ($M$) which decays according to the spatial and temporal distance between earthquakes. Then, the intensity function in the ETAS model is given by
$\lambda^{*}_{\mu,\nu}(t,x,y|H_{t})=\mu(x,y)+\sum_{\{i:t_{i}<t\}}\nu(t-t_{i},x-x_{i},y-y_{i},M_{i})$ (1)
where $(x_{i},y_{i})$ are the coordinates of the epicenter and $t_{i}$ the time of occurrence of the event $i$, with i=1,2,,Ni=1,2,...,N. Physically, $\mu(x,y)$ denotes the background seismicity, which refers to the earthquake rate that is unaffected by the occurrence of any other seismic event. The triggering function $\nu$ is described as the decay of an earthquake’s ability to generate aftershocks over time (i.e., it is based on the Omori Law [utsu1995centenary]). Then, only earthquakes with magnitude exceeding an umbral $M_{u}$ are considered, based on the magnitude completeness provided by the Gutenberg-Richter Law[gutenberg1949seismicity]. Then, the log-likelihood of the model is given by
$\ell(\mu,\nu):=\sum_{k=1}^{n}\log(\lambda^{*}_{\mu,\nu}(t_{k},x_{k},y_{k}|H_{t_{k}}))-\int_{0}^{T}\iint_{S}\lambda^{*}_{\mu,\nu}(t,x,y|H_{t})dxdydt,$ (2)
As we mentioned in Section 1, the catalog from [sawires2019updated] allows us to focus solely on the Poissonian intensity function. Therefore, in the rest of the text, we will work with a Non-homogeneous Poisson Process (NHPP) with intensity function defined as:
$\lambda(t,x,y)=\mu(x,y,t),$ (3)
to model the background seismic function.
It is important to note that if we have a sample $X=\{(x_{i},y_{i},t_{i})\}$, where $(x_{i},y_{i})$ are observed in a region $D\subset\mathbb{R}^{2}$, with i=1,2,,Ni=1,2,...,N and in a time interval $(0,T)$ from a NHPP the likelihood [daley2003introduction] is given by
$L(X|\mu)=e^{\iint_{D}\int_{0}^{T}\mu(x,y,t)dtdxdy}\prod_{i=1}^{N}\mu(x_{i},y_{i},t_{i}),$ (4)
which, in general, is enormously computationally expensive to compute because it requires solving a triple integral. Consequently, authors such as [kottas2007bayesian, zhuang2002stochastic, ross2022semiparametric], among others, have assumed, in the stationary case, that $\mu(x,y)$ is a bounded function. This form allows modeling it as
$\mu(x,y)=\gamma f(x,y)$
where $\gamma$ is in $\mathbb{R}^{+}$, and $f(x,y)$ is a density function. Additionally, if we assume that the domain of $f(x,y)$ is $D$, then the likelihood for the homogeneous case (in time) equation (4) can be rewritten as
$L(X|\mu)=e^{-\gamma T}\gamma^{N}\prod_{i=1}^{N}f(x_{i},y_{i}),$
which significantly reduces the computational cost, and then the idea of using the Dirichlet process to model $\mu$ as in [kottas2007bayesian] or [ross2022semiparametric] for a Bayesian approach emerges.
As we discussed in 1 following the idea of writing $\mu(x,y)=\gamma f(x,y)$, authors as [li2024multivariate] and [xiao2015modeling] have proposed defining $\mu$ piecewise constant in time, i.e., if we have a partition of the interval (0,T) given by times $0=S_{1},S_{2},...,S_{p+1}=T$, then we can define $\mu$ as
$\mu(x,y,t)=\sum_{p=1}^{P}\gamma_{p}f_{p}(x,y)\mathbbm{1}(S_{p}<t<S_{p+1})$ (5)
where $\mathbbm{1}(\cdot)$ is the indicator function, which is the definition of $\mu$ that is used in our work. This definition aims to work with the inhomogeneities over space and time, but also allows us to rewrite the equation (4) as
$L(X|\mu)=\prod_{p=1}^{P}\Big(e^{-\gamma_{p}|I_{p}|}\gamma^{n_{p}}\prod_{i=1}^{n_{p}}f_{p}(x_{p,i},y_{p,i})\Big),$ (6)
where $I_{p}=(S_{p},S_{p+1})$, $|I_{p}|$ is its longitude, $X_{p}=\{(x_{p,i},y_{p,i})\}$ with $i=1,...,n_{p}$ is the subsample of $X$ of elements where $t_{i}$ is in $I_{p}$ for p=1,,Pp=1,...,P.
As can be seen in this definition of $\mu$, we have created non-exchangeable subsamples of $X$ that are dependent on $\prod_{i=1}^{n_{p}}f_{p}(x_{p,i},y_{p,i})$, which is a reasonable model if dependency is assumed. The temporal dependency used in this work can be described as a particular case of GDP, which directed acyclic graph is shown in Figure 1.
The GDP [chakrabarti2024graphical] allows working with a more general DAG. Nevertheless, for our purposes, this version is adequate, and it also coincides with structures previously studied in the nested Chinese restaurant process [blei2010nested] or the Nested Hierarchical Dirichlet Processes [paisley2014nested] that are englobed in the GDP framework.
[…]
Figure 1: Graphical Dirichlet Process denoting time dependency.
As mentioned by [muller2015bayesian], and since the Dirichlet processes are discrete with probability one, it is awkward to estimate continuous densities. This trouble could be fixed using DPM, which is generalized in [chakrabarti2024graphical] to the GDP. They proved that the structure presented in figure 1 used to define a GDP mixture is
$G_{1}|\alpha_{1},G_{0}$ $\sim DP(\alpha_{1},G_{0})$
$G_{p}|\alpha_{p},G_{p-1}$ $\sim DP(\alpha_{p},G_{p-1})$
and for each group $p$ we have
$\theta_{p,l}|G_{p}$ $\stackrel{{\scriptstyle\text{ind}}}{{\sim}}G_{p}$
$(x_{p,i},y_{p,i})|\{\theta_{p,l}\}_{l=1}^{L}$ $\sim F(\{\theta_{p,l}\}_{l=1}^{L})$
which can be recovered as the limit of a hierarchical finite model when the number of $L$ mixture elements tends to infinity. Then, the addition of $\gamma_{p}$ variables, a priori independent to all $G_{p}$ processes in our hierarchical model, is given by
$\alpha_{1}|\alpha_{0}$ $\sim\text{Gamma}(\alpha_{0},1)$ (7)
$\beta_{1}|\alpha_{1}$ $\sim\text{Dir}\Big(\frac{\alpha_{1}}{L},...,\frac{\alpha_{1}}{L}\Big)$
$\alpha_{t}|\alpha_{t-1}$ $\sim\text{Gamma}(\alpha_{t-1},1)$
$\beta_{p}|\alpha_{p},\beta_{p-1}$ $\sim\text{Dir}(\alpha_{p}(\beta_{p-1,1},\beta_{p-1,2},...,\beta_{p-1,L}))$
$\psi_{l}|G_{0}$ $\sim G_{0}$
$z_{pi}|\beta_{t}$ $\sim\text{Cat}(1:L,\beta_{p})$
$\gamma_{p}$ $\sim\text{Gamma}(\gamma_{0}k,k)$
$\{(x_{p,i},y_{p,i})|z_{p,i},\{\psi_{l}\}_{l=1}^{L},\gamma_{p}\}_{i=1}^{n_{p}}$ $\sim\mathcal{F}(z_{p,i},\{\psi_{l}\}_{l=1}^{L},\gamma_{p})\text{ for all $p$ in \{1,2,..,P\}}$
where $\{(x_{p,i},y_{p,i})|z_{p,i},\{\psi_{l}\}_{l=1}^{L},\gamma_{p}\}_{i=1}^{n_{t}}\sim\mathcal{F}(z_{p,i},\{\psi_{l}\}_{l=1}^{L},\gamma_{t})$ has the following probability density function
$f_{x}((x_{p,1},y_{p,1}),...,(x_{p,n_{p}},y_{p,n_{p}}))=e^{-\gamma_{p}|I_{p}|}\gamma_{t}^{n_{p}}\prod_{i=1}^{n_{p}}\Big(\prod_{l=1}^{L}\phi_{\psi_{l}}(x_{p,i})^{\mathbbm{1}(z_{pi}=l)}\Big).$
In this work, we assume that the functions $\phi_{\mu_{l},\Sigma^{-1}_{l}}$, l=1,2,,Ll=1,2,...,L are Gaussian probability densities with mean $\mu_{l}$ and covariance matrix $\Sigma_{l}$. Also, to facilitate the notation, we write $\psi_{l}=(\mu_{l},\Sigma_{l}^{-1})$ and $\phi_{\psi_{l}}$. Although this could break the assumption that $f$ has support $D$, we assume, as in [ross2022semiparametric], that the probability outside of $D$ is negligible and therefore Equation (6) is still valid.
The main reason for assuming that $f_{\psi_{l}}$ is Gaussian is to obtain the full conditional of $\psi_{l}$ using conjugate analysis, which will allow its incorporation into a Markov Chain Monte Carlo (MCMC) straightforwardly.
It is worth mentioning that [kottas2007bayesian] proposed a bivariate beta distribution to guarantee bounded support in $D$. Nevertheless, based on the previous paragraph, we decided to follow the approach proposed by [ross2022semiparametric].
For all our examples, we define $G_{0}\sim\text{NIW}(\mu_{0},\eta,\Sigma_{0},\nu)$ where NIW denotes a Normal Inverse Wishart distribution, also we denote his density by $f_{\text{NIW}}(\cdot)$. Therefore, our posterior distribution is given by
$f(\boldsymbol{\alpha},\boldsymbol{\beta},\boldsymbol{\gamma},\boldsymbol{\psi},\mathbf{Z}|\boldsymbol{Y})$ $\propto\alpha_{1}^{\alpha_{0}-1}e^{-\alpha_{1}}\frac{\displaystyle\prod_{l=1}^{L}\beta_{1l}^{\frac{\alpha_{1}}{L}-1}}{\text{B}(\frac{\alpha_{1}}{L},\frac{\alpha_{1}}{L},...,\frac{\alpha_{1}}{L})}\Big(\prod_{p=1}^{P-1}\frac{\alpha_{p+1}^{\alpha_{p}-1}e^{-\alpha_{p+1}}}{\Gamma(\alpha_{p})}\Big)$
$\Big(\prod_{p=2}^{P}\frac{\displaystyle\prod_{l=1}^{L}\beta_{pl}^{\alpha_{p}\beta_{p-1,l}-1}}{\text{B}(\alpha_{p}\beta_{p-1,1},...,\alpha_{p}\beta_{p-1,L})}\Big)\Big(\prod_{l=1}^{L}f_{\text{NIW}}(\psi_{l})\Big)$
$\Big(\prod_{p=1}^{P}\prod_{i=1}^{n_{p}}\prod_{l=1}^{L}\beta_{pl}^{\mathbbm{1}(z_{pi=l})}\Big)\Big(\prod_{p=1}^{P}\gamma_{p}^{\gamma_{0}k-1}e^{-\gamma_{p}k}\Big)$
$\prod_{p=1}^{P}e^{-\gamma_{p}|I_{p}|}\gamma_{p}^{n_{p}}\prod_{i=1}^{n_{p}}\Big(\prod_{l=1}^{L}\phi_{\psi_{l}}(x_{p,i},y_{p,i})^{\mathbbm{1}(z_{pi}=l)}\Big)$
where
$\boldsymbol{\alpha}$ $=(\alpha_{1},...,\alpha_{p})$
$\boldsymbol{\gamma}$ $=(\gamma_{1},...,\gamma_{p})$
$\boldsymbol{\psi}$ $=(\psi_{1},...,\psi_{L})$
$\boldsymbol{Z}$ $=(z_{1,1},...,z_{1,n_{1}},z_{2,1},...,z_{2,n_{2}},...,z_{p,n_{p}})$
$\boldsymbol{\beta}$ $=\begin{pmatrix}\beta_{11}&\ldots&\beta_{1L}\\ \vdots&\ddots&\vdots\\ \beta_{p1}&\ldots&\beta_{pL}=\end{pmatrix}=\begin{pmatrix}\beta_{1}\\ \vdots\\ \beta_{p}\end{pmatrix}$
where the full conditional for $\gamma_{p}$ for all $p$ in 1,…,$P$ is distributed as $\text{Gamma}(\gamma_{0}k+n_{p},k+|I_{p}|)$, for $\psi_{l}$ with $l=1,..,L$ we have
$\pi(\psi_{l}|\cdot)\propto\prod_{p=1}^{P}\prod_{i=1}^{n_{p}}\Big(\prod_{l=1}^{L}\phi_{\psi_{l}}(x_{p,i},y_{p,i})^{\mathbbm{1}(z_{pi}=l)}\Big)f_{\text{NIW}}(\psi_{l})$
which by conjugated analysis is distributed as $\text{NIW}(\mu_{l},\eta_{l},\Sigma_{l},\nu_{l})$ where
$\mu_{l}$ $=\frac{\eta\mu_{0}+m_{l}\bar{y}}{\eta+m_{l}}$
$\eta_{l}$ $=\eta+m_{l}$
$\nu_{n}$ $=\nu+m_{l}$
$\Sigma_{l}$ $=\Sigma_{0}+S+\frac{\eta m_{l}}{\eta+m_{l}}(\bar{y}-\mu_{0})(\bar{y}-\mu_{0})$
defined by
$m_{l}$ $=\displaystyle\sum_{p=1}^{P}\sum_{i=1}^{n_{p}}\mathbbm{1}(z_{p,i}=l)$
$\bar{y}$ $=\frac{\displaystyle\sum_{p=1}^{P}\sum_{i=1}^{n_{p}}(x_{p,i},y_{p,i})\mathbbm{1}(z_{p,i}=l)}{m_{l}}$
$S$ $=\displaystyle\sum_{p=1}^{P}\sum_{i=1}^{n_{t}}((x_{p,i},y_{p,i})-\bar{y})((x_{p,i},y_{p,i})-\bar{y})^{T}\mathbbm{1}(z_{p,i}=l),$
in the case of $\boldsymbol{\alpha}$ we have
$\pi(\alpha_{1}|\cdot)$ $\propto\frac{e^{-\alpha_{1}}\alpha_{1}^{\alpha_{0}}\alpha_{2}^{\alpha_{1}}}{\Gamma(\alpha_{1})}\frac{\displaystyle\prod_{l=1}^{L}\beta_{1l}^{\frac{\alpha_{1}}{L}}}{\text{B}(\frac{\alpha_{1}}{L},\frac{\alpha_{1}}{L},...,\frac{\alpha_{1}}{L})}$
$\pi(\alpha_{P}|\cdot)$ $\propto e^{-\alpha_{P}}\alpha_{P}^{\alpha_{P-1}}\frac{\displaystyle\prod_{l=1}^{L}\beta_{Pl}^{\alpha_{P}\beta_{P-1,l}}}{\text{B}(\alpha_{P}\beta_{P-1,1},...,\alpha_{P}\beta_{P-1,L})}$
and for all $p$ in 2,…,P1P-1
$\pi(\alpha_{p}|\cdot)$ $\propto\frac{e^{-\alpha_{p}}\alpha_{p}^{\alpha_{p-1}}\alpha_{p+1}^{\alpha_{p}}}{\Gamma(\alpha_{p})}\frac{\displaystyle\prod_{l=1}^{L}\beta_{pl}^{\alpha_{p}\beta_{p-1,l}}}{\text{B}(\alpha_{p}\beta_{p-1,1},...,\alpha_{p}\beta_{p-1,L})}$
also, for $\boldsymbol{\beta}$ we have
$\pi(\beta_{1}|\cdot)\propto\prod_{l=1}^{L}\beta_{1l}^{m_{1l}+\frac{\alpha_{1}}{L}-1}\frac{\displaystyle\prod_{l=1}^{L}\beta_{2l}^{\alpha_{2}\beta_{1,l}}}{\text{B}(\alpha_{2}\beta_{1,1},...,\alpha_{2}\beta_{1,L})}$
meanwhile for $p$ in 2,,P2,...,P we have $\beta_{p}|\cdot\sim\text{Beta}(m_{pl}+\alpha_{p}\beta_{pl},...,m_{pl}+\alpha_{p}\beta_{pl})$ where $m_{pl}=\displaystyle\sum_{i=1}^{n_{p}}\mathbbm{1}(z_{p,i}=l)$, finally
$\mathbb{P}(Z_{pi}=l|\cdot)\propto\beta_{pl}\phi_{\psi_{l}}(x_{p,i},y_{p,i})$
[…]
As discussed in [chakrabarti2024graphical], the total conditional distributions of $\alpha_{i}$, i=1,,Pi=1,...,P, and $\beta_{1}$ are not standard distributions and are the main bottleneck in the MCMC. Since all $\alpha_{i}$ are positive random variables, we proposed a random walk movement to explore them. According to [chakrabarti2024graphical], we used proposals based on the SALTSampler introduced by [director2017efficient] to explore $\beta_{1}$. Then, we opted for a Hybrid MCMC [robert1999monte] to sample the posterior distribution; the code is available at github.com/isaiasmanuel/NHPP.
Since the discussion by Veen and Schoenberg [veen2008estimation], the use of hidden variables has increased in popularity to estimate the ETAS model. This is because, as they pointed out, the direct optimization of the likelihood, as originally proposed by Zhuang et al. [zhuang2002stochastic], is computationally expensive and numerically unstable. The hidden variables included are
$\chi_{ii}$ $=\begin{cases}1,&\text{if earthquake $i$ is a background event}\\ 0,&\text{otherwise}\end{cases}$ (8)
$\chi_{ij}$ $=\begin{cases}1,&\text{if earthquake $i$ is an aftershock of $j$}\\ 0,&\text{otherwise}\end{cases},$ (9)
where s=0,,ms=0,...,m. These variables represent the branch structure of the earthquakes, and the log-likelihood of the ETAS model incorporating them is given by
$\ell_{c}\Big(\mu,\eta,\{\chi_{ii}\}_{i=1}^{N},\{\chi_{ij}\}_{(i,j)\in I^{2}}\Big)=\ell^{*}_{\text{O}}(\mu,\{\chi_{ii}\}_{i=1}^{N})+\ell^{*}_{\text{I}}(\nu,\{\chi_{ij}\}_{(i,j)\in I^{2}}\ ),$
where $I^{2}$ denotes $\{1,2,...,N\}\times\{1,2,...,N\}$, being $\times$ the cartesian product, $\ell_{\text{O}}(\cdot)$ is the log-likelihood due to the offspring, and $\ell_{\text{I}}(\cdot)$ is the log-likelihood due to immigrants. They are given by
$\ell^{*}_{\text{I}}(\mu,\{\chi_{ii}\}_{i=1}^{N})$ $=\sum_{i=1}^{n}\chi_{ii}\log(\mu(x_{i},y_{i}))-\int_{0}^{T}\iint_{S}\mu(x,y,t)dxdydt$ (10)
$\ell^{*}_{\text{O}}(\nu,\{\chi_{ij}\}_{(i,j)\in I^{2}}\ )$ $=\sum_{j=1}^{n}\Big[\sum_{i>j}\chi_{ij}\log\Big(\nu(t_{i}-t_{j},x_{i}-x_{j},y_{i}-y_{j},M_{j})\Big)$ (11)
$\hskip 42.67912pt-\int_{t_{j}}^{T}\iint_{S}\nu(t-t_{j},x-x_{j},y-y_{j},M_{j})dxdydt\Big].$
Given this framework, [veen2008estimation, li2020space] have used the EM algorithm for the ETAS model, and then [molkenthin2022gp, ross2022semiparametric] have introduced Bayesian hierarchical schemes where Gibbs proposals are used. It is important to highlight that only the Poissonian term of the model presented in this section can be easily incorporated into the ETAS framework by redefining $\ell^{*}_{\text{I}}$ as
$\ell^{*}_{\text{I}}(\mu,\{\chi_{ii}\}_{i=1}^{N})=\sum_{p=1}^{P}\Big(\sum_{i=1}^{n}\chi_{ii}\log(\mu^{p}(x_{i},y_{i},t_{i}))-\int_{0}^{T}\iint_{S}\mu^{p}(x,y,t)dxdydt\Big).$
However, in our case, this evaluation is not necessary to estimate the background intensity function because the features of the [sawires2019updated] catalog allow us to simply assume that all the $\chi_{ij}$ are 0 for all $i\neq j$, whereas $\chi_{ii}=1$ for all $i$.
3 Numerical Experiments
3.1 Simulated Data
To verify the performance of our model, we apply it to a synthetic earthquake catalog, with $D=(-5,10)\times(-5,10)$ and $t\in(0,10)$ with the following intensity function
$\lambda(x,y,t)=\Big(50\mathbbm{1}(t<5)+100\mathbbm{1}(t\geq 5)\Big)\Big(h(t)g_{1}(x,y)+(1-h(t))g_{2}(x,y)\Big)$ (12)
where $\mathbbm{1}(\cdot)$ denotes the indicator function and
$h(t)$ $=\frac{1}{1+e^{-\frac{t-T}{2}}}$
$g_{1}(x,y)$ $=\frac{2}{3}\phi_{(0,0),\mathbf{1}_{2\times 2}}(x,y)+\frac{1}{3}\phi_{(2,2),\mathbf{1}_{2\times 2}}(x,y)$
$g_{2}(x,y)$ $=\frac{2}{3}\phi_{(6,2),\mathbf{1}_{2\times 2}}(x,y)+\frac{1}{3}\phi_{(4,6),\mathbf{1}_{2\times 2}}(x,y)$
The idea behind the proposed $\lambda$ function is to create a spatio-temporal density function that can capture the typical challenges observed in an earthquake catalog. For example, the number of recorded earthquakes may increase due to changes in instrumentation or aseismic transients as we discussed in Section 1.
In this example we take for the Normal-inverse-Wishart distribution $\mu_{0}=(1,1)$, $\Sigma_{0}=\mathbb{I}_{2}$, $\nu=3$ and $\eta=0.1$, where $\mathbb{I}_{2}$ is the $2\times 2$ identity matrix, the idea is to have a non informative prior distribution for the parameters in the mixture. For all the $\gamma_{p}$ we take $\gamma_{0}=70$ and $k=0.1$, which results in a vague prior distribution and $\alpha_{0}=1$. Also, a regular partition of the interval $(0,T)$ was used with $P=8$ and $L=8$.
Figure 2: $\lambda$ evaluated at different times
Figure 2 shows the intensity function in the whole space at different times. This visual representation facilitates the observation of spatial and temporal changes. The realization of this NHPP was generated using the thinning algorithm [lewis1979simulation2, illian2008statistical]. The simulated catalog and the code to generate it can be consulted in github.com/isaiasmanuel/NHPP.
Figure 3: Posterior mean of $\lambda$
Figure 4: Posterior mean of $\lambda$ with transparency based on the coefficient of variation using simulated data
Figure 3 presents the posterior mean of $\gamma_{p}f_{p}$, as defined in Equation (5). Then, to utilize the information in the posterior distribution more effectively, we present the Figure 4. The colors of the figure are the same as in Figure 3, but now we include a transparency based on the posterior coefficient of variation (CV), which is defined by
$\text{CV}_{p}(x,y):=\frac{\sigma(x,y)}{\mu(x,y)},$
where $\mu$ and $\sigma$ are the mean and posterior variance of $\lambda(x,y,t)$ where $t\in(S_{p},S_{p+1})$. Then the transparency for each figure $p$ is given by
$1-\frac{\text{CV}_{p}-\min_{x,y,p}\text{CV}_{p}}{\text{CV}_{p}},$
In other words, if the posterior mean is high compared to the posterior standard deviation, the color is saturated. The transparency increases as the average decreases relative to the standard deviation.
As can be seen in Figure 4, we have lower coefficient of variation values in regions with higher values of $\lambda$ (i.e., and as expected, the regions with the highest number of points are better estimated). To complement Figure 4, we also present Figure 5, where the histograms of the marginal posterior samples are presented for each $\gamma_{p}$ and the real value $\gamma_{p}$ (vertical dashed blue line), according to Equation (12). Figure 5 reveals that the posterior distributions of $\gamma_{p}$ adequately contain the real value for all the $p=1,2,$....
Figure 5: Histogram of the posterior $\gamma_{p}$, p=1,2,,Pp=1,2,...,P using simulated data. Blue dashed lines are the real $\gamma_{p}$ for each interval.
Figure 6: Clustering of the simulated data using 4 groups.
In this work, we do not focus on the inference related to each element in the mixture. Instead, we are interested in inferring the intensity function $\lambda$. Nevertheless, it is worth mentioning that we observed the label switching problem, as expected in a finite mixture problem [green1995reversible, zhang2004learning, bouguila2012fully]. For this reason, we have implemented the algorithm proposed by [papastamoulis2010artificial] (the code is available at github.com/isaiasmanuel/NHPP), mainly to select the thinning in the chain.
Finally, to validate our model in Figure 6, we present the clustering of the observed points. We show the mean of the Gaussian distributions (red crosses) used to define $\lambda$ in Equation 12. As can be seen, we recover the four mixture components adequately, even when the mixture weights change over time. The clusters were taken using the least-square criterion proposed in [dahl2006model], as our affinity matrix for spectral clustering [von2007tutorial], and the number of clusters was taken using the eigengap heuristic.
From this example, we have seen how GDP can be used to accurately estimate the intensity function of an inhomogeneous spatial Poisson process. That is why, in the next section, we apply this same methodology to the [sawires2019updated] catalog to recover the background seismic function.
3.2 Southern Mexico
As we discussed previously in Section 1, different authors have pointed out that the background seismicity function could change over time by aseismic transients [matthews1988statistical, marsan2013slow, reverso2015detection, nishikawa2023development] or by strong earthquakes [melgar2018deep, cruz2021short]; Nevertheless, a common assumption in the ETAS model is to assume a constant background seismicity function [zhuang2002stochastic, ross2022semiparametric, molkenthin2022gp], if this assumption is suitable for the seismicity in southern Mexico, we expect to see homogeneity between all $\gamma_{p}f_{p}$ for the whole p=1,2,,Pp=1,2,...,P. Alternatively, if the assumption is not appropriate, we expect to observe changes over space or time related to the regions where we observed SSE or Strong earthquakes.
As reported by [radiguet2012slow, radiguet2016triggering, cruz2021short], the SSE in Guerrero, Mexico, between 2000/01/01 and 2015/12/31, had an approximate periodicity of 4 years; for this reason, in the GDP inference for the Mexican data, we use a regular time partition with $P=4$, thus obtaining an SSE at each time interval.
The epicenters taken from [sawires2019updated] are presented in Figure 7 (colored by their time interval $p$). The Middle America Trench [GEBCO] is presented in fuchsia. Furthermore, the slip contour curves digitized from [radiguet2012slow] of the 2001-2002, 2006, and 2009-2010 SSE are presented in blue, orange, and red, respectively (each 4 cm). In addition, the 15 cm slip contour of the 2014 SSE is presented in black, which was digitized from [radiguet2016triggering]. Additionally, we present the slip distribution, with 1-meter resolution, for the earthquakes occurred in $2003/01/22$, $2012/03/20$, and $2014/04/18$, with magnitudes 7.6, 7.4 and 7.2, and epcenters -104.1040E° 18.770N°, -98.2310E° 16.493N°, -100.9723E° 17.397N°, respectively, using data from [Geological].
Figure 7: Epicenters from [sawires2019updated] catalog, SSE from [radiguet2012slow, radiguet2016triggering], slip distribution contours for earthquakes with $M\geq 7$ from [Geological] and the Middle America Trench using data from [GEBCO].
The hyperparameters in the Normal Inverse Wishart were taken as $\nu=6$, $\eta=0.01$, $\Sigma_{0}=2\mathbb{I}_{2}$, $\mu_{0}=(-102,17)$ and we restrict $\mu$ to be in the observed domain (i.e., $(-105.5\text{E}^{\degree},-96.5\text{N}^{\degree})\times(15\text{E}^{\degree},19.5\text{N}^{\degree})$). In this case, the prior is vague for $\mu_{l}$ with l=1,2,,Ll=1,2,...,L inside the observed domain, and we are being informative for $\Sigma_{l}$, in the sense that we have
$\mathbb{E}((\Sigma_{l})_{11})=\frac{2}{3},\quad\text{Var}((\Sigma_{l})_{11})=\frac{8}{9},$
To analyze the sensitivity of our model to this prior, we compared the results obtained by performing inference using the non-informative prior from the previous example. The distribution of $\gamma_{p}f_{p}$ did not show any notable changes, but the mixing velocity of the MCMC increased significantly using the new proposal.
We take $\alpha_{0}=1$ as in the previous example. For $\gamma_{p}$, we follow [ross2022semiparametric], and we take $\gamma_{0}=1$ and $k=0.01$. Finally, we set $L=12$, similar to the paper by [chakrabarti2024graphical], where $L=10$ was used, and a simulation example with a similar-sized sample compared to ours was approximated. However, we ran MCMC with a larger $L$ value, but we did not observe improvements in the estimated intensities. In contrast, the MCMC worsened due to the label switching, as there were constantly mixture components with 0 observations allocated.
Figure 8: Posterior mean of $\lambda$ in $\frac{\text{events}}{\text{days deg${}^{2}$}}$ with transparency based on the coefficient of variation using the catalog of [sawires2019updated]. The gray curves in $p=1,2,3$ are the slip contours each 4 cm for the corresponding SSE in Figure 7 and in $p=4$ is the slip contour at 15 cm.
Figure 8 shows the posterior mean of $\gamma_{p}f_{p}$ for each $p=1,2,3,4$. It can be observed that the intensity function reaches its maximum near the slip distribution contours of the strong earthquakes in the interval $p=1$ and $p=4$. Furthermore, we can also appreciate that near the border of the SSE slips, the background seismic function exhibits significant activity. Previously, [fukuda2018variability] have observed a similar behavior in the Boso Peninsula, Japan. The increase in seismic activity at the border of the SSE could be explained by stress redistribution caused by the largest SSE slips, which may reduce stress in some regions while increasing it in others, potentially triggering earthquakes where stress accumulated but was not released, as previously hypothesized in the region by [cruz2021short, cruz2025seafloor, villafuerte2025slow], among others.
Figure 9: Histogram of the posterior $\gamma_{p}$, p=1,2,,Pp=1,2,...,P using the catalog of [sawires2019updated].
While Figure 8 shows the spatio-temporal changes, Figure 9 presents the marginal posterior distributions of $\gamma_{p}$ for $p=1,2,\ldots,4$. In particular, we observe a significant shift toward higher values in the $\gamma_{4}$ distribution, indicating an increase in the expected number of earthquakes during the last period, which is consistent with the occurrence of the two strong earthquakes observed in this interval.
Figure 10: Clustering using the catalog of [sawires2019updated] using 8 groups, each color represents one group, and the geometric figures represent the period of time.
As we presented in Section 3.1, one of the main advantages of using a Dirichlet process mixture is that we are also clustering the events[muller2015bayesian] when estimating $f_{p}$. In Figure 10, we present the data from [sawires2019updated] clustered in 8 groups using the [dahl2006model] methodology.
[…]
4 Conclusions
Although the use of DPM for non-parametric Bayesian estimation of the background seismic activity function had been previously studied by [ross2022semiparametric], they only considered homogeneity in the time case. In this work, we have succeeded in considering temporal changes over time in the background activity function by including GDP [chakrabarti2024graphical].
Furthermore, in the context of estimating the background seismic activity function with GDP, it has allowed us to easily identify seismogenic regions by only taking seismicity into account, which had not been previously discussed by [ross2022semiparametric] and which is not intuitive to obtain following the approach presented in [molkenthin2022gp].
As discussed previously by [molkenthin2022gp], it is important to note that the Gaussian process approach requires the correlation structure to be specified, whereas the DP approach requires the number of mixture components to be determined. The use of a mixture of DPs can estimate a more flexible dependence structure, mainly due to local behaviors, than Gaussian processes.
Our work can be viewed as an extension of the ideas presented in [matthews1988statistical], which proposed using a non-homogeneous Poisson process for the background activity function. The intensities estimated using the catalog by [sawires2019updated] for southern Mexico have exhibited spatial and temporal changes associated with strong regular earthquakes and SSEs. Therefore, we can conclude that the assumption of a homogeneous Poisson process is not realistic for southern Mexico.